Please help me solve this
If the length of a side of a square is increased by 3 centimeters, the perimeter of the new square is 40 centimeters more than twice the length of a side of the original square. Find the dimensions of the original square

Answer by stanbon(75887)

If the length of a side of a square is increased by 3 centimeters, the perimeter of the new square is 40 centimeters more than twice the length of a side of the original square. Find the dimensions of the original square
-----
General Formula: P = 4s where s is a side of the square.
===============
Your Problem:
P = 4s
2s+40 = 4(s+3)
2s+40 = 4s+12
2s = 28
---
s = 14 cm
---
Dimensions of the original square: 14 cm by 14 cm
============================
Cheers,
Stan H.